OpenStudy (anonymous):

can some one explain this to me let f(x)=x squared -81. FInd f1(x)

OpenStudy (accessdenied):

Do you mean to find \(f^{-1} (x) \) ?

OpenStudy (anonymous):

no that is what the question says thats why im confused

OpenStudy (anonymous):

\[f(x)=x ^{2}-81 find f ^{1}(x)\]

OpenStudy (accessdenied):

What topic are you working on in Math? Maybe it is \(f'(x) \), as in the derivative of f(x) with respect to x?

OpenStudy (anonymous):

functions

OpenStudy (accessdenied):

Hm... I do not know what that notation is supposed to mean, then. I want to say it is a typo for inverse function of f. Are there answer choices with this problem as well?

OpenStudy (anonymous):

yes the answer choices are

OpenStudy (anonymous):

\[\pm \sqrt{x+81}\]

OpenStudy (anonymous):

\[9\sqrt{x}\]

OpenStudy (anonymous):

\[1/x ^{2}-81\]

OpenStudy (anonymous):

\[x ^{2}/81\]

OpenStudy (accessdenied):

Alright. I definitely believe this is meant to be \(f^{-1} (x) \) of \(f(x) = x^2 - 81\) now, as I can see the answer among the choices. You know how to find inverse functions?

OpenStudy (anonymous):

i don't now how to do it when there is a number behind the \[x ^{2}\]

OpenStudy (accessdenied):

Have you seen this method? We switch \(f(x)\) for x, and x for \(f^{-1}(x) \) in the equation. \(x = \left( \color{blue}{f^{-1}(x)} \right)^2 - 81 \) Where we just have to solve for the function like we would any variable: \(x = \color{blue}{y}^2 - 81 \)

OpenStudy (anonymous):

i have never seen the first method but have seen the second method

OpenStudy (accessdenied):

They're basically the same, but I'll continue writing with y. \( x = y^2 - 81 \) We first solve for the y^2. We can do that by adding 81 to both sides, cancelling the -81 next to y^2. Then take the positive and negative square root of both sides. That would get y by itself with no powers.

OpenStudy (anonymous):

how would you take to square root of x+81

OpenStudy (accessdenied):

We wouldn't need to simplify it, we just leave it like this: \( \pm \sqrt{x + 81} = y \) We can take the square root, even if the left side doesn't necessarily simplify nicely like square root of 9 = 3. Like if we had 6 = y^2, we could take + or - sqrt(6) = y without having to simplify 6. It is the same with x + 81. :)

OpenStudy (anonymous):

oh ok so wouldn't that be the answer

OpenStudy (accessdenied):

Yes, that should be the answer. :)

OpenStudy (anonymous):

ok thank you so much

OpenStudy (accessdenied):

Glad to help!